# Sparse Mixture of Experts as Unified Competitive Learning

### v1

@@ Abstract
Sparse Mixture of Experts (SMoE) improves the efficiency of large language model training by directing input tokens to a subset of experts. Despite its success in generation tasks, its generalization ability remains an open question. In this paper, we demonstrate that current SMoEs, which fall into two categories: (1) Token Choice ;and (2) Expert Choice, struggle with tasks such as the Massive Text Embedding Benchmark (MTEB). By analyzing their mechanism through the lens of competitive learning, our study finds that the Token Choice approach may overly focus on irrelevant experts, while the Expert Choice approach risks discarding important tokens, potentially affecting performance. Motivated by this analysis, we propose Unified Competitive Learning SMoE (USMoE), a novel and efficient framework designed to improve the performance of existing SMoEs in both scenarios: with and without training. Extensive experiments across various tasks show that USMoE achieves up to a 10% improvement over traditional approaches or reduces computational inference costs by 14% while maintaining strong performance.
Sparse Mixture of Experts as Unified Competitive Learning
Giang Dothanks: Corresponding author  Hung Le  Truyen Tran Applied Artificial Intelligence Institute (A2I2), Deakin University {s224363215,thai.le,truyen.tran}@deakin.edu.au
@@ 1 Introduction
Sparse Mixture of Experts (SMoE) models have achieved notable success in natural language processing (NLP) and visual representation learning tasks (Du et al., 2022; Fedus et al., 2022; Riquelme et al., 2021a; Shen et al., 2023). These advancements build on the Transformer architecture (Vaswani et al., 2017) and its variants (Child et al., 2019; Dai et al., 2019b), which leverage large datasets and significant compute resources. However, training large Transformer models can be prohibitively expensive, requiring extensive compute hours (Kaddour et al., 2023). To overcome this issue, SMoE models activate only a subset of experts for each input, reducing inference time compared to dense models (Shazeer et al., 2017a; Zoph et al., 2022; Artetxe et al., 2022; Krajewski et al., 2024). The SMoE architecture can be categorized into two variants: Token Choice, which assigns experts to each token Dai et al. (2024); Team (2024); Muennighoff et al. (2024); Jiang et al. (2024a), and Expert Choice, which assigns tokens to each expert Zhou et al. (2022b). The advantage of Token Choice lies in its ability to dynamically select experts for each token, while Expert Choice ensures a more balanced token distribution across experts.
Figure 1: We compare the performance of USMoE (ours) with the Expert Choice and Token Choice approaches on the Massive Text Embedding Benchmark (MTEB). The results demonstrate that our method outperforms traditional approaches across six tasks using OLMoE-7B  Muennighoff et al. (2024) without additional training.
@@
Despite their promising results, SMoE models have several limitations. The Expert Choice approach suffers from token dropping  Zhou et al. (2022b), while the Token Choice approach struggles with unbalanced expert loading  Shazeer et al. (2017b). Additionally, both approaches are prone to representation collapse, where either a few experts dominate the routing or all experts learn similar representations Chi et al. (2022a); Chen et al. (2022). Recent research has explored improving router policies Chi et al. (2022b); Chen et al. (2023a); Do et al. (2023a) to mitigate these issues. However, existing methods face two key challenges: (1) the use of auxiliary losses requires balancing router loss and task loss, leading to trade-offs, and (2) they still struggle with the fundamental limitations of either the Token Choice or Expert Choice approach.
Which approach is better for SMoE - Token Choice or Expert Choice - in terms of generalization across multiple tasks, both with and without training?
In this paper, we address this question by reexamining SMoE through the lens of Competitive Learning Rumelhart and Zipser (1985); Kaski and Kohonen (1994); Srivastava et al. (2013); Pham et al. (2024a). From this perspective, Token Choice can be seen as horizontal competitive learning, where the most similar expert is selected for each token, while Expert Choice represents vertical competitive learning, where each expert selects the most similar tokens. This viewpoint reveals a key trade-off: Expert Choice risks dropping important tokens, whereas Token Choice must process both relevant and irrelevant tokens.
@@
Building on this analysis, we propose Unified Competitive Learning SMoE (USMoE), a robust and efficient framework comprising two key components: (1) Unified Competitive Score and (2) Unified Competitive Mechanism. These components enable the SMoE model to dynamically prioritize tokens or experts while ensuring the selection of the most similar token-expert pair, enhancing both robustness and effectiveness. To demonstrate the effectiveness of our approach, we evaluate USMoE across multiple scenarios, including pretraining and both fine-tuned and non-fine-tuned settings. USMoE consistently outperforms baseline methods across these scenarios, with particularly strong gains in tasks that require deep input understanding, such as semantic textual similarity, classification, and clustering. Extensive experiments across various benchmarks show that USMoE achieves up to a 10% improvement over traditional approaches or reduces inference computational costs by 14%, all while maintaining high performance.
@@
  • We introduce a Competitive Learning perspective on SMoE, highlighting the weaknesses of existing approaches.
  • We propose USMoE, a robust and efficient framework that addresses the limitations of both Token Choice and Expert Choice.
@@
  • We conduct extensive experiments on large language models, covering pretraining and both fine-tuned and non-fine-tuned settings, providing a detailed analysis of USMoE’s performance and effectiveness.
@@ 2 Related Work
Sparse Mixture of Experts (SMoE). Sparse Mixture of Experts (SMoE), an extension of the Mixture of Experts framework (Jacobs et al., 1991; Jordan and Jacobs, 1994), has gained traction with large language models and has since been applied in various domains, including computer vision and speech recognition (Zhou et al., 2022c; Riquelme et al., 2021b). The SMoE architecture consists of two main variants: Token Choice, where experts are assigned to each token (Shazeer et al., 2017b; Fedus et al., 2022; Jiang et al., 2024b; Do et al., 2024b), and Expert Choice, where tokens are assigned to specific experts (Zhou et al., 2022b).
Token Choice treats all tokens equally, which has raised concerns among researchers (Wu et al., 2021; Hou et al., 2022; Lin et al., 2025), while Expert Choice suffers from token-dropping issues. Additionally, SMoE faces the challenge of representation collapse, where experts produce similar outputs. Various solutions have been proposed, such as XMoE, which employs low-dimensional routing scores (Chi et al., 2022b), and SMoE-dropout, which gradually activates more experts (Chen et al., 2023a). Other approaches, including HyperRouter (Do et al., 2023a) and StableMoE (Dai et al., 2022a), focus on enhancing router stability and robustness. Although these advancements have improved SMoE models, representation collapse remains a persistent issue (Pham et al., 2024a; Do et al., 2024a). Our approach addresses this by optimizing the alignment between tokens and the most suitable experts, expanding expert specialization and mitigating collapse.
Competitive Learning. Competitive learning is an unsupervised learning approach where computational units compete to respond to a given input, enabling feature discovery (McClelland and Rumelhart, 1987; Andersen et al., 1969; Stefanis and Jasper, 1969). Inspired by the biological brain, this concept has recently been leveraged to enhance the efficiency of Large Language Models (LLMs) (Cai et al., 2024; Zhao et al., 2024). Furthermore, competitive learning has been shown to enhance the efficiency of Mixture of Experts (MoE) models (Ahn and Sentis, 2021; Cai et al., 2023).
The most relevant work to our work is CompeteSMoE (Pham et al., 2024b), which frames competition through expert outputs and employs a router trained to predict competition outcomes in a scheduled manner. However, it shares a key limitation with traditional competitive learning frameworks, activating all experts, which becomes infeasible for large-scale models with millions of parameters. In contrast, we redefine competition among expert embeddings into two categories: Token Choice SMoE, which represents horizontal competitive learning, and Expert Choice SMoE, which represents vertical competitive learning. Building on this perspective, we introduce two novel scoring methods and mechanisms - Unified Competitive Score and Unified Competitive Mechanism - that enhance traditional SMoE in both performance and efficiency.
@@ 3 Methodology
We introduce Unified Competitive Learning SMoE (USMoE), a novel and efficient Sparse Mixture of Experts framework designed to address the limitations of both Token Choice and Expert Choice through a unified competitive learning mechanism. As shown in Section 3.2, our approach consists of two key components:
  • Unified Competitive Score – a scoring function that balances expert selection.
  • Unified Competitive Mechanism – a structured routing strategy that ensures efficient and effective expert allocation.
@@ 3.1 Preliminaries
Sparse Mixture of Experts.  The Sparse Mixture of Experts (SMoE) is typically a transformer architecture that replaces the multi-layer perceptron (MLP) layers in standard transformers with Mixture of Experts (MoE) layers, inspired by (Shazeer et al., 2017a). Given $\mathbf{x}\in\mathbb{R}^{b\times d}$ as the output of the multi-head attention (MHA) layer, the result of the SMoE with $n$ experts is a weighted sum of each expert’s computation, $E_{i}(x)$, weighted by the router function $\mathcal{S}(x)$:
$f_{\mathrm{SMoE}}(\boldsymbol{x})=\sum_{i=1}^{n}\mathcal{S}(\boldsymbol{x})_{i}\cdot E_{i}(\boldsymbol{x})$ (1)
Where $\mathcal{S}(x)$ is computed by $TopK$ function as below the Equation  2, and $W_{e}$ is a learnable experts embeddings.
$\mathcal{S}(\boldsymbol{x})=\operatorname{TopK}(\operatorname{softmax}(W_{e}\cdot x),k)$ (2)
$\operatorname{TopK}(\boldsymbol{v},k)=\begin{cases}\boldsymbol{v_{i}}&\text{if}\boldsymbol{v_{i}}\text{ is in the top }k\text{ largest of }\boldsymbol{v},\\ -\infty&\text{otherwise}.\end{cases}$
3.2 Unified Competitive Learning SMoE (USMoE)
Unified Competitive Score.  As illustrated in Figure 2, the Token Choice approach selects and promotes the best expert by comparing the similarity scores between a token and the available experts. Given a representation $\mathbf{x}\in\mathbb{R}^{b\times l\times d}$, which is the output of the multi-head attention (MHA) layer, and a router matrix $W_{e}\in\mathbb{R}^{d\times n}$, where $n$ represents the number of experts, the similarity score $s_{t}$ in the Token Choice approach is computed as:
$s_{t}=\operatorname{softmax}(W_{e}\cdot x,d=-1)$ (3)
where $d$ denotes the dimension along which the softmax function is applied. In contrast, the Expert Choice approach selects and promotes the best tokens by comparing the similarity scores between an expert and the available tokens. The similarity score $s_{e}$ in the Expert Choice approach is computed as:
$s_{e}=\operatorname{softmax}(W_{e}\cdot x,d=1)$ (4)
While Expert Choice ensures an equal distribution of tokens across experts, Token Choice guarantees that each token is fairly processed by the model. To leverage the strengths of both approaches, we introduce the Unified Competitive Score, a soft combination of the Token Choice and Expert Choice scores, defined as:
$s_{u}=\alpha\cdot s_{e}+(1-\alpha)\cdot s_{t}$ (5)
where $\alpha\in[0,1]$ is a tunable control parameter that can be adjusted based on the data. In practice, we find that $\alpha\approx 0.5$ is an appropriate choice.
Unified Competitive Mechanism.  Token Choice selects experts using horizontal competitive learning, where each token is assigned to the most similar expert. This approach performs well when there is a highly relevant expert for a given token. However, its effectiveness diminishes when the similarity scores between the token and all experts are low. In contrast, Expert Choice selects tokens using vertical competitive learning, where each expert is assigned the most similar tokens. This approach is effective when the similarity score distribution across experts is well-structured. However, this condition is difficult to achieve due to the issue of representation collapse  Chi et al. (2022b). As a result, Expert Choice may either discard important tokens or lead to multiple experts selecting the same token. USMoE addresses these challenges by introducing the Unified Competitive Mechanism, which treats expert and token selection as a joint (expert, token) pairing process. As demonstrated in Algorithm  1, we first flatten the similarity matrix and then select the top-$N$ highest-scoring pairs, representing the most similar expert-token combinations.
Let the ordered array $(S^{t}_{1},S^{t}_{2},\dots,S^{t}_{N})$ represent the top-$N$ token scores selected using the Token Choice approach, where $S^{t}_{1}\leq S^{t}_{2}\leq\dots\leq S^{t}_{N}$. Similarly, let $(S^{e}_{1},S^{e}_{2},\dots,S^{e}_{N})$ denote the top-$N$ token scores selected using the Expert Choice approach, where $S^{e}_{1}\leq S^{e}_{2}\leq\dots\leq S^{e}_{N}$. Finally, let $(S^{u}_{1},S^{u}_{2},\dots,S^{u}_{N})$ represent the top-$N$ token scores selected using the Unified Competitive Score approach, where $S^{u}_{1}\leq S^{u}_{2}\leq\dots\leq S^{u}_{n}$. We derive the following inequalities, with a detailed proof provided in the Appendix A.1:
$S^{t}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N]$ (6)
$S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N]$
Based on Inequality 6, the Unified Competitive Mechanism ensures that the selected expert-token pairs are at least as optimal as those chosen by the Token Choice approach. Moreover, this mechanism addresses the limitations of Expert Choice, which may drop important tokens, and Token Choice, which may select irrelevant tokens.
Figure 2: An illustration of our USMoE framework, which enhances existing SMoE models through a Unified Competitive Learning approach. The method first flattens the score matching between tokens and experts into a 1D representation, then selects the Top-N best pairs, and finally maps the results back to the original 2D matching scores. Best viewed in color.

### v2

# Unified Sparse Mixture of Experts

@@ Abstract
Sparse Mixture of Experts (SMoEs) models scale the capacity of models while maintaining constant computational overhead. Early designs typically relied on a fixed value of $k$, where $k$ represents either the number of experts selected per token or the number of tokens assigned per expert. However, these approaches encounter three key limitations: they may fail to route to important experts or tokens, may assign irrelevant ones, and often suffer from representation collapse among experts. This paper reexamines SMoEs through the lens of Linear Programming, and proposes a Unified Sparse Mixture of Experts (USMoE) framework that addresses these limitations. Specifically, our approach introduces a unified mechanism that integrates information from both the expert and token dimensions, and a unified scoring function that linearly combines similarity scores between experts and tokens. We provide both theoretical justification and empirical evidence demonstrating USMoE’s effectiveness in overcoming the limitations of traditional routing methods. Through comprehensive evaluations on both clean and corrupted settings for large language models and vision tasks, under both training-free and training scenarios, USMoE achieves up to a 10% performance improvement over standard approaches or reduces inference costs by up to 14%, while maintaining competitive accuracy.
@@ 1 Introduction
Sparse Mixture of Experts (SMoE) models have achieved notable success in natural language processing (NLP) and visual representation learning tasks (du_glam_2022, ; fedus_switch_2022, ; riquelme2021scalingvisionsparsemixture, ; shen-etal-2023-scaling, ). These advancements build on the Transformer architecture (NIPS2017_3f5ee243, ) and its variants (child2019generating, ; dai2019transformerxl, ), which leverage large datasets and significant compute resources. However, training large Transformer models can be prohibitively expensive, requiring extensive compute hours (kaddour2023challenges, ). To overcome this issue, SMoE models activate only a subset of experts for each input, reducing inference time compared to dense models (shazeer2017outrageously, ; zoph2022stmoe, ; artetxe2022efficient, ; krajewski2024scaling, ). The SMoE architecture can be categorized into two variants: Token Choice, which assigns experts to each token dai2024deepseekmoeultimateexpertspecialization ; qwen_moe ; muennighoff2024olmoeopenmixtureofexpertslanguage ; jiang2024mixtral , and Expert Choice, which assigns tokens to each expert zhou_mixture_experts_2022 . The advantage of Token Choice lies in its ability to dynamically select experts for each token, while Expert Choice ensures a more balanced token distribution across experts.
Input
USMoE
TC
EC
Input
USMoE
TC
EC
Figure 1: We compare token routing performance in vision tasks using 7×7 images, where each token is color-coded based on its assigned expert under setting $c=t$, where $c$ is a sparsity constraint and $t$ is number of image patches. Left: When the object is easy to distinguish, Expert Choice (EC) fails to assign different experts appropriately. Token Choice (TC) performs better but still does not align perfectly with the actual object, while USMoE correctly separates the object. Right: In more challenging images, both Expert Choice (EC) and Token Choice (TC) fail to distinguish between object and background. In contrast, USMoE successfully differentiates the object from the background, demonstrating greater efficiency in vision tasks compared to EC and TC, as further shown in Section 4.
@@
Despite their promising results, SMoE models have several limitations. The Expert Choice approach suffers from token dropping  zhou_mixture_experts_2022 , while the Token Choice approach struggles with unbalanced expert loading  shazeer2017outrageouslylargeneuralnetworks . Additionally, both approaches are prone to representation collapse, where either a few experts dominate the routing or all experts learn similar representations chi_representation_2022 ; chen_towards_2022 . Recent research has explored improving router policies chi2022representation ; chen2023sparse ; do2023hyperrouter to mitigate these issues. However, existing methods face three key challenges: (1) The reliance on auxiliary losses requires careful balancing between the router loss and the task loss, which introduces trade-offs; (2) Token Choice (TC) struggles to handle noisy tokens effectively; and (3) Expert Choice (EC) suffers from information leakage issues, which significantly degrade performance on autoregressive models zhou2022mixtureofexpertsexpertchoicerouting ; wang2024auxiliarylossfreeloadbalancingstrategy ; raposo2024mixtureofdepthsdynamicallyallocatingcompute . As a result, the question of how to optimally select experts or tokens remains open.
Research Question What are the optimal criteria for selecting experts or tokens in SMoE models under fixed computational budgets?
Loading Balance Robustness Token Dropping Information Leak Expert Collapses
Token Choice Φ Φ Φ
Expert Choice Φ Φ
USMoE (ours) Φ Φ Φ Φ Φ
Table 1: Benefits of the proposed USMoEs’ model form over existing methods.
In this paper, we revisit SMoEs through the lens of Linear Programming NIPS2015_09662890 ; NEURIPS2021_a8fbbd3b to better understand the criteria for expert and token selection. From this perspective, Token Choice selects experts along the expert dimension by assigning each token to its most similar expert, while Expert Choice selects tokens along the token dimension by allowing each expert to choose the most similar tokens.
This perspective highlights a key trade-off: Expert Choice risks dropping important tokens, whereas Token Choice struggles to handle noisy or irrelevant tokens. Furthermore, both approaches suffer from the problem of representation collapse chi2022representation ; do2023hyperrouter ; pham2024competesmoe .
@@
Building on this analysis, we propose the Unified Sparse Mixture of Experts (USMoE), a robust and efficient framework consisting of two key components: (1) the Unified Score and (2) the Unified Mechanism. The Unified Score is defined as a linear combination of two mapping functions that transform the dot product similarity between token and expert embeddings into a probability distribution - using functions such as softmax or sigmoid, as shown in Algorithm 1. In parallel, the Unified Mechanism incorporates information from both the token and expert dimensions, as illustrated in Figure 2. These components enable the SMoE model to dynamically prioritize tokens or experts while ensuring the selection of the most similar token-expert pair, enhancing both robustness and effectiveness. We summarize the benefits of USMoEs’ model form over existing methods in Table  1. To demonstrate the effectiveness of our approach, we evaluate USMoE across a variety of scenarios, including training-free, pretraining, and fine-tuning settings, spanning both language and vision domains under both clean and corrupted datasets. USMoE consistently outperforms baseline methods across these scenarios, with particularly strong gains in tasks that require deep input understanding, such as semantic textual similarity, classification, and clustering. Extensive experiments across various benchmarks show that USMoE achieves up to a 10% improvement over traditional approaches or reduces inference computational costs by 14%, all while maintaining high performance.
@@
  • We introduce a Linear Programming perspective on SMoE, highlighting the weaknesses of existing approaches.
  • We propose USMoE, a robust and efficient framework comprising two modules: (1) the Unified Score and (2) the Unified Mechanism, which together address the limitations of conventional approaches, as summarized in Table 1.
@@
  • We conduct extensive experiments on both large language models and vision domains, covering pretraining and both fine-tuning and training-free settings, providing a detailed analysis of USMoE’s performance and effectiveness.
@@ 2 Related Work
Sparse Mixture of Experts (SMoE). Sparse Mixture of Experts (SMoE), an extension of the Mixture of Experts framework (jacobs1991, ; jordan1994, ), has gained traction with large language models and has since been applied in various domains, including computer vision and speech recognition (NEURIPS2022_2f00ecd7, ; NEURIPS2021_48237d9f, ). The SMoE architecture consists of two main variants: Token Choice, where experts are assigned to each token (shazeer2017outrageouslylargeneuralnetworks, ; fedus_switch_2022, ; jiang2024mixtralexperts, ; do2024effectivenessdiscreterepresentationssparse, ), and Expert Choice, where tokens are assigned to specific experts (zhou_mixture_experts_2022, ).
Token Choice treats all tokens equally, which has raised concerns among researchers (wu2021attentionneed, ; hou2022tokendroppingefficientbert, ; lin2025rho1tokensneed, ), while Expert Choice suffers from token-dropping issues. Additionally, SMoE faces the challenge of representation collapse, where experts produce similar outputs. Various solutions have been proposed, such as XMoE, which employs low-dimensional routing scores (chi2022representation, ), and SMoE-dropout, which gradually activates more experts (chen2023sparse, ). Other approaches, including HyperRouter (do2023hyperrouter, ) and StableMoE (dai2022stablemoe, ), focus on enhancing router stability and robustness. Although these advancements have improved SMoE models, representation collapse remains a persistent issue (pham2024competesmoe, ; do2024simsmoesolvingrepresentationalcollapse, ). Our approach addresses this by optimizing the alignment between tokens and the most suitable experts, expanding expert specialization and mitigating collapse. Moreover, we also discuss related work on Linear Programming and other studies relevant to our approach in Section A.1.
@@ 3 Methodology
We introduce Unified Sparse Mixture of Experts (USMoE), a novel and efficient Sparse Mixture of Experts framework designed to address the limitations of both Token Choice and Expert Choice through a Unified Score, a scoring function that balances expert and token important and Unified Mechanism, a structured routing strategy that consider information from both token dimension and expert dimension.
@@ 3.1 Preliminaries
The Sparse Mixture of Experts (SMoE) architecture replaces the MLP layers in standard transformers with Mixture of Experts (MoE) layers (shazeer2017outrageously, ). Let $h\in\mathbb{R}^{T\times d}$ denote the token representations from the attention layer, where $T$ is the number of tokens and $d$ is the hidden dimension. Given $N$ expert functions $\{\text{FFN}_{j}\}_{j=1}^{N}$, a gating projection matrix $W\in\mathbb{R}^{d\times N}$ maps tokens to expert affinity scores. Denotes $S=f(hW)\in\mathbb{R}^{T\times N}$ is a compatibility score, where $f$ maps the scores to a routing distribution (e.g., softmax or sigmoid). Given $X\in\{0,1\}^{T\times N}$ be a binary routing matrix, where $x_{ij}=1$ if token $i$ is routed to expert $j$, the output of the SMoE layer is:
$f_{\text{SMoE}}(h)=\sum_{j=1}^{N}S_{\cdot j}\odot X_{\cdot j}\odot\text{FFN}_{j}(h\odot X_{\cdot j}),$ (1)
where $\text{FFN}_{j}(\cdot)$ denotes the computation performed by expert $j^{th}$, and $\odot$ represents element-wise multiplication, and FFN denotes for the Feed-forward neural networks.
3.2 Experts Selection as Linear Programming
Viewing expert selection as a linear programming problem allows for globally optimizing expert assignments by maximizing similarity scores under a strict sparsity (computational) constraint.
Expert Selection Optimization.  Given a computational budget of $c$ experts per token, the objective of Sparse Mixture-of-Experts (MoE) routing is to determine the optimal expert assignment $X^{*}$ that maximizes the total compatibility score, subject to a sparsity constraint $c$:
maximize $\sum_{i=1}^{T}\sum_{j=1}^{N}S_{ij}x_{ij}$ (2)
subject to $\sum_{i=1}^{T}\sum_{j=1}^{N}x_{ij}\leq c$
$x_{ij}\in\{0,1\},\quad\forall i,j.$
Figure 2: An illustration of our USMoE selection mechanism (middle) is shown, which incorporates information from both the token and expert dimensions. In contrast, Token Choice (TC, left) considers only the expert dimension, while Expert Choice (EC, right) focuses solely on the token dimension. Tokens marked with # represent noisy or irrelevant tokens. TC struggles to handle these # tokens effectively, and EC is prone to missing important tokens. USMoE addresses both issues, making our method more robust than traditional MoE approaches. This robustness is demonstrated in both theoretical analysis and experimental results. ’E’ denotes experts. Best viewed in color.
Proposition 3.1.
Let $S\in\mathbb{R}^{T\times N}$ be the compatibility score matrix and $c\in\mathbb{N}$ a global routing budget. Consider the objective:
$M=\langle S,X\rangle=\sum_{i=1}^{T}\sum_{j=1}^{N}S_{ij}x_{ij},$
subject to the constraint $\sum_{i,j}x_{ij}\leq c$, with $x_{ij}\in\{0,1\}$.
Let $X_{\text{USMoE}}=\text{TopK}(S,c)$ be the binary mask produced by selecting the top-$c$ entries of $S$ globally. Then for any other feasible binary routing matrix $X_{T}\in\{0,1\}^{T\times N}$ satisfying the same budget constraint $\sum_{i,j}(X_{T})_{ij}\leq c$, we have:
$\langle S,X_{T}\rangle\leq\langle S,X_{\text{USMoE}}\rangle.$
The Proposition 3.1 indicate that the Unified Mechanism (USMoE) yields the optimal solution that maximizes the total similarity under the global budget constraint. The proof is provided in Appendix  A.
Definition 3.2 (Unified Score Function).
Let $S\in\mathbb{R}^{T\times N}$ be the compatibility score matrix between $T$ tokens and $N$ experts. Define two score mapping functions:
  • $f_{t}(S)\in\mathbb{R}^{T\times N}$: a row-wise scoring function used in Token Choice (e.g., softmax applied across each row).
  • $f_{e}(S)\in\mathbb{R}^{T\times N}$: a column-wise scoring function used in Expert Choice (e.g., softmax applied across each column).
The Unified Score Function combines both perspectives via a linear combination:
$f_{\text{USMoE}}(S)=\alpha\cdot f_{e}(S)+\beta\cdot f_{t}(S),$
where $\alpha,\beta\in\mathbb{R}$ are non-negative coefficients such that $\alpha+\beta=1$.
The Unified Score as Definition 3.2 integrates both token-centric and expert-centric preferences to inform more balanced routing decisions.
Lemma 3.3.
The Unified Score Function $f_{\text{USMoE}}(S)=\alpha\cdot f_{e}(S)+\beta\cdot f_{t}(S)$, where $\alpha,\beta\in\mathbb{R}$ are non-negative coefficients such that $\alpha+\beta=1$, is more robust to representation collapse chi2022representation than using $f_{t}(S)$ or $f_{e}(S)$ alone.
Lemma 3.3 establishes that USMoE outperforms conventional SMoEs by mitigating the representation collapse issue, as discussed in chi2022representation . A detailed proof is presented in Appendix A.
3.3 Unified Sparse Mixture of Experts (USMoE)
Revising Expert Choice. The Expert Choice (EC) approach faces scalability challenges in autoregressive models due to information leakage wang2024auxiliarylossfreeloadbalancingstrategy ; raposo2024mixtureofdepthsdynamicallyallocatingcompute introduced by the softmax operator. To address this, we replace softmax with a sigmoid function for Token Choice, which eliminates inter-token leakage and, as shown in Figure 6, leads to improved bits-per-character (BPC) performance and a reduced token dropping ratio.
Finding 1: The Sigmoid function outperforms Softmax in Expert Choice MoE, as it handles information leakage and token dropping more effectively.
Unified Score. We empirically observe that Token Choice effectively captures in-context learning behavior, while Expert Choice conveys rich semantic alignment as Figure 3. To harness the complementary strengths of both, we propose the Unified Score, defined in Algorithm 1, as a weighted sum of the Token Choice and Expert Choice scores. This unified formulation integrates both token-centric and expert-centric signals to enable more balanced and informed routing decisions. Empirically, we find that setting the combination weight to $\alpha\approx 0.5$ yields a robust trade-off between the two strategies.
Figure 3: The performance of Unified Score (USMoE), Softmax (TC), and Sigmoid (EC) across the MTEB benchmark. Sigmoid outperforms Softmax on tasks without prompting, indicating stronger semantic representations.
Finding 2: Sigmoid (EC) excels at capturing semantic information, while Softmax (TC) is better suited for leveraging in-context cues.
Unified Mechanism. We propose the Unified Mechanism, which formulates expert-token assignment as a joint selection problem. By flattening the similarity matrix and selecting the top-$N$ expert-token pairs based on the highest unified scores, as illustrated in Figure 3.1, this mechanism enables efficient, context-aware routing within sparse MoE architectures.
[…]
2:Output $Y$
[…]
8:return $Y$